Technical Observing - 20140207 UT

Observers: JHill (Tucson - remote room)
Telescope Operator: SAllanson (LBTO)
AO Support:: (Tucson)
Telescope Support:: MMidkiff (LBTO)
Instrument Support:: JPower (LBTO)
Engineering Support: (Tucson)

Summary

This technical night is a mix of E-time (TCS tests and instrument configuration tests) and D-time (staff science and AZ science). The weather is mostly cloudy with snow expected. We never opened.

The patches to MCSPU, GCS and ECSGUI all worked without any problems, so they are qualified to remain in place if needed.

Plan

  • DONE Verify that moving systems are all functional after this morning's unplanned power outage.

  • DONE Test new version of MCSPU r752

  • DONE GCS patch 2.14.17

  • DONE prototype ECSGUI 3.0 (/home/mdelapena/ECSGUI/ECSGUI )

  • DONE Follow the chronic problems ...... we'll put some more mileage on RFBG rotator and DX primary cell hardware.

  • DONE We'll also check out some of the binocular collimation adjustments from Issue 5030.

  • DONE We'll run some MODS1/LBCred observations in symbiotic mode (Issue 4513).

Details

Moving Systems Checkout

See Steve's notes in the night log .....

Doug, Juan Carlos and Marco have checked out both AdSecs

Olga has checked MODS1, and Michelle has checked LBCs.

ECSGUI 3.0

01:30 UT In terms of display of parameters, ECSGUI 3.0 is apparently identical with the old ECSGUI - even down to the font mistakes. We are testing it in terms of command button functionality as the night progresses. LGT controls are OK. MV controls work. VD controls work.

GCS 2.14.17

The link to the executable was changed this afternoon, and both GCSs have been restarted.

02:00 Mirror ventilation running as primaries are 4 degC above outside ambient.

02:05 Side vent doors open midway.

Marco is power cycling DX AdSec - Crate #5 is not responding properly.

Authorizing as MODS1 on SX only.

02:08 Steve turns on azcamserver.

Steve inits hexapods and then M3s.

02:18 Fake guide stars in GCS cp -p gcs.conf.2sides.overlay.Star gcs.conf
Steve restarts GCSs.

02:22 TRACK preset for MODS1

SX M1 X=0.074 Y=0.037 Z=-0.211 RX=-7.37 RY=19.2
SX M2 X=-2.242 Y=+1.173 Z=2.572 RX=294.9 RY=-53.9

02:27 Marco reports DX M2 is now OK. Perhaps something in the fastlink communication needed to warm up? The dome was about -2 deg all day while the unit was powered off.

02:27 GUIDE preset with MODS1 succeeds with fake guide stars.

Offset succeeds.

02:31 ACTIVE preset with MODS1

Binocular Collimation

02:35 Authorize as MODS1 on left and LUCI2 on right

02:29 SYNC ACTIVE presets at EL=68

SX M1 X=-1.031 Y=-0.010 Z=-0.221 RX=1.076 RY=-52.5
SX M2 X=-6.950 Y=+0.526 Z=2.653 RX=304.5 RY=-123.5
DX M1 X=+0.658 Y=+0.893 Z=0.000 RX=-50.648 RY=52.64
DX M2 X=+2.547 Y=-3.898 Z=3.297 RX=-129.9 RY=-72.0

For the moment, we decide to do nothing, and monocular SX is the best range-balanced configuration.

RFBG Rotator

Is turned on since ~1 UT, even though we aren't particularly using it until 02:30.

02:55 Authorize as LUCI2 on RFBG only

02:57 ACTIVE preset for LUCI

DX M1 X=-0.700 Y=+1.018 Z=0.000 RX=-59.1 RY=-39.2
DX M2 X=-3.612 Y=-3.048 Z=3.336 RX=-135.5 RY=-165.7

Pseudo Monocular Tests - Closed Dome

03:02 Authorize MODS1 on left in PMM, and LBC-Red on the right

03:03 Async preset to MODS1

DX M1 X=-1.409 Y=-0.435 Z=-0.357 RX=-21.9 RY=+84.4
SX M1 X=-0.477 Y=-0.187 Z=0.230 RX=18.4 RY=-12.5
SX M2 X=-4.236 Y=-0.688 Z=2.737 RX=323.8 RY=-80.1

03:22 Turn on LBC

03:30 Running tests of pseudo-monocular mode. Script pmm_observe.py has been improved to cancel the OB if it sees the telescope go off-source while the OB is running.

04:00 We are now closed for humidity as well as heavy clouds.

02/07 4:32 lbcobs@obs2 ~/supportscripts/SymbioticSupport$ ./pmm_observe.py -s DX -t 100.0 -r z_RedSymbiotic.ob 
./pmm_observe.py script version of 07-February-2014
Passive Side Requested is  right   Scaling Exposure Time by  100.0
Red OB file: z_RedSymbiotic.ob
Looping to watch for telescope on-source -- ^C to exit

05:15 Ice has accumulated on the all-sky camera. We officially have no chance of opening.

05:18 fake guide stars have been removed.

05:19 Telescope moving to zenith

05:20 LBC turned off

05:22 MODS1 put to sleep

-- JohnHill - 07 Feb 2014
Topic revision: r2 - 09 Feb 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback